Peanuts cost $4 per pound and walnuts cost $5.50 per pound. A 40-pound mixture sells for $184. How many pounds of each kind of nut are in the mixture?
Answer by sofiyac(983)


Let x be the number of pounds of peanuts, then the number of pounds of walnuts will be 40-x.
The 40 pound mixture sells for 184, so 184/40=4.6 dollars per pound of the final product.
4x+5.50(40-x)=4.6*40
4x+220-5.50x=184
-1.50x+220=184 subtract 220 from each side
-1.50x=184-220
-1.50x=-36 divide both sides by -1.50
x=24
So 24 pounds of peanuts and 40-24=16 pounds of walnuts.